solve triangle ABC given that a=13 b=14 and c=11. round to the nearest tenth.
i hope you can help
thanks a bunch!

Answer by stanbon(75887)

solve triangle ABC given that a=13 b=14 and c=11. round to the nearest tenth.
Use the Law of Cosines to get:
CosA = (b^2+c^2-a^2)/(2bc)
CosA = 0.48
A=arccos(0.48)=61.31 degrees
-----------
CosB = (a^2+c^2-b^2)/(2ac)
CosB = 0.3287
B=arccos(0.3287)=70.81 degrees
---------------
C=180-(61.31+70.81)=47.88 degrees
---------
Cheers,
Stan H.